Please solve it by hand or typing, don't use EXCEL.
For D Draw the time line to show cash flow.

1. XYZ stock price and dividends history are as follows Year Beginning-of-year price Dividend paid at year- $100 $110 $90 $95 end $4 $4 $4 $4 2005 2006 2007 2008 An investor buys three shares of XYZ at the beginning of 2005, buys another two shares at the beginning of 2006, sells one share at the beginning of 2007, and sells all four remaining shares at the beginning of 2008. Hint: When the investor buys three shares at the beginning of 2005, she cannot receive dividends paid at the end of 2004. She can receive dividend payments on these three shares at the end of 2005 A. Computer HPR for each period (05-06, 06-07, and 07-08) (6pts) B. Compute the arithmetic return. (3pts) C. Compute the geometric return. (3pts) D. Computer the dollar-weighted rate of return. Draw the time line to show cash flows. Write down the formula and calculate the dollar-weighted rate of return. The time line includes four dates corresponding to the turns of the year for January 1, 2005 to January 1, 2008. (8pts)
Answer #1

At the end of the year appreciation is below :

2005: 330-300+3*4=42

2006: 450-550+ 5*4=80

2007:380-360+4*4=36

Return is calculated as appreciation over the initial value of portfolio:

2005: 42/300 =14%

2006:80/550 =14.545%

2007:36/260=14%

Arthematic average of the numbers above (14%--14.545%+10%)/3=3.15%

Geometric average=-1+[(1+0.14)*(1-0.14545)*(1+0.1)]^(1/3) = 2.33%

For Dollar return --0.17% you need to use excel or ffinancial calculator you get the cash flow .see below.

cash flow

2005:-300

2006: -208

2007: 110

2008: 396

Now use = IRR() function the answer will be -0.17%